the sum of the areas of two circles is 612pi square meters. Find the length of a radius of the smaller circle if one of them is 4 times as long as the other.
Answer by mananth(16946)

let the smaller circle radius be r m
larger circle radius = 4r m
...
area of smaller circle +area of larger circle = 612 pi
pi*r^2 + pi*(4r)^2= 612*pi
pi(r^2+16r^2)=612*pi
/pi
17r^2= 612
/17
r^2= 612/17
r^2= 36
r=6 m the radius of smaller circle.
...
m.ananth@hotmail.ca